OpenStudy (mpj4):

how to solve this equation? sqrt{ (x-1)^2 + (y+2)^2 } = sqrt{ (x-5)^2 + (y-4)^2 } + 1

OpenStudy (mpj4):

\[\sqrt{\left ( x-1 \right )^{2}+\left ( y+2 \right )^{2}}=\sqrt{\left ( x-5 \right )^{2}+\left ( y-4 \right )^{2}}+1\]

OpenStudy (mpj4):

ayy lmao. just figured it out xD

OpenStudy (anonymous):

How did you do it? Square twice (shifting terms around after the first time), then deal with the plus/minus stuff? Or something nicer?